Problem 4 Intro Lomack Companys bonds have a 11-year maturity, a 10% coupon, paid semiannually, and a par value of $1,000. The market interest rate is 3%, with semiannual compounding. Part 1 What is the bonds price (in $)? B Attempt 1/10 for 10 pts. No decimals Submit
Problem 5 Intro A corporate bond has 16 years to maturity, a face value of $1,000, a coupon rate of 4.9% and pays interest twice a year. The annual market interest rate for similar bonds is 3.1%. Part 1 What is the price of the bond (in $)? -Attempt 1 /10 for 10 pts. No decimals Submit
B Attempt 1/10 for 10 pts. Part2 2 years later, the market interest rate for similar bonds has gone up to 4.1 %, what is the new price of the bond (in $)? No decimals Submit
Answer #1

4.1). To find the bond's price, we need to put the following values in the financial calculator:

INPUT 11x2=22 3/2=1.5 (10%/2)x1,000=50 1,000
TVM N I/Y PV PMT FV
OUPUT -1,651.73

So, Bond's Price = $1,651.73

5.1). To find the bond's price, we need to put the following values in the financial calculator:

INPUT 16x2=32 3.1/2=1.55 (4.9%/2)x1,000=24.5 1,000
TVM N I/Y PV PMT FV
OUPUT -1,225.71

So, Bond's Price = $1,225.71

5.2). To find the bond's price, we need to put the following values in the financial calculator:

INPUT 14x2=28 4.1/2=2.05 (4.9%/2)x1,000=24.5 1,000
TVM N I/Y PV PMT FV
OUPUT -1,084.58

So, Bond's Price = $1,084.58